OpenStudy (anonymous):

find the equation of the circle with the center (-2,9) and radius of 3 plz show work

OpenStudy (anonymous):

(x+2)^2+(y-9)^2=9

OpenStudy (anonymous):

just plug what you were given into the standard equation of a circle: (X-H)^2 + (Y-K)^2=r^2

OpenStudy (anonymous):

so te radius is 3?

OpenStudy (anonymous):

the radius

OpenStudy (anonymous):

you said the radius is 3 in the question.

OpenStudy (anonymous):

it is three in the question.. i guess im confused

OpenStudy (anonymous):

look at the equation of the circle, what happens to the radius when we plug it in, radius is represented by r

OpenStudy (anonymous):

center \[(h,k)\] radius \[r\] circle is \[(x-h)^2+(y-k)^2=r^2\] as lagrange said

OpenStudy (anonymous):

thnks guys